OpenStudy (mathhater):

Give the 5th and 10th term for the sequence formula below. tn = (n – 3)2

OpenStudy (callisto):

I assume the general term to be \[t_n = (n-3)^2\] \[t_5 = (5-3)^2 = 2^2 = 4\] Can you try the 10th term?

OpenStudy (mathhater):

can you show me how to do it please?

OpenStudy (callisto):

Just put n = 10 into the general term and simplify it. Like I put n = 5 into \(t_n = (n-3)^2 \) and simplify it